OpenStudy (anonymous):

needing portfolio help 10th grade algebra quadratics unit 4 portfolio i got the answer but my teacher said i didn't answer it the quadratics way.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

don't you hate it when you do it right but not the teacher way? what the monkey is "quadratic way" supposed to be? probably an outtake from a Van Morrison album your teacher listened to in 1978

OpenStudy (anonymous):

haha i guess it wasnt the way i was suppose to do it but the answer i got in the end was right just i didnt show the work the way it was suppose to be shown but i did show my work so i didnt get credit so i have to redo it i just dont understand what i do now.

OpenStudy (freckles):

can you post the question

OpenStudy (anonymous):

m: A ball is dropped from the top of a building that is 250 feet tall. The height h of the ball in feet after t seconds is modeled by the function h = -16t2 +250. How long will it take for the ball to reach the ground? Round your answer to the nearest tenth if necessary.

OpenStudy (anonymous):

Interpret the concepts of the task and translate them into mathematics. (RESTATE THE PROBLEM IN YOUR OWN WORDS AND HOW WILL YOU USE MATH TO SOLVE THE PROBLEM)

OpenStudy (freckles):

they want to know when h=0

OpenStudy (anonymous):

State your answer clearly. Notice that the correct answer is only 25% of your grade!

OpenStudy (freckles):

so you solve 0=-16t^2+250

OpenStudy (anonymous):

): State the solution/outcome in the context of the task. Defend the process, evaluate and interpret the reasonableness of the solution/outcome. (CHECK YOUR ANSWER)

OpenStudy (freckles):

and how did you solve that exactly so we know not to do that way

OpenStudy (anonymous):

this is what i got Distance= one half acceleration times time times time (1/2 at^2) To solve for time distance must equal the height the ball falls from. (250 feet)

OpenStudy (anonymous):

250=32t^2 15.625=t^2 3.95=t ( nearest 10th= 4 seconds)

OpenStudy (misty1212):

i guess you should say that when it hits the ground the height is zero (doe) so your task is to solve \[-16t^2+250=0\] for \(t\) this should take exactly two steps

OpenStudy (misty1212):

ok 3 steps \[250=16t^2\\ \frac{250}{16}=t^2\\ \sqrt{\frac{250}{16}}=t\] you can clean that up a bit

OpenStudy (anonymous):

but i dont understand what way she wants me to solve it?

OpenStudy (anonymous):

@misty1212